# DurFlex-EVC: Duration-Flexible Emotional Voice Conversion with Parallel Generation

### v3

@@ III-C Unit Aligner
The unit aligner is introduced to model the semantic context in which frame-level features are compressed to the unit level, which works similar to an information bottleneck. Fig. 1d shows the structure of the unit aligner. We combined learnable embeddings with cross-attention to derive attention weights for specific embeddings a feature focuses on, which is used for duration modeling. The attention weights are trained to predict the unit through the index of the largest value, and the duration is modeled as the consecutive number of units obtained. We use the output of the style autoencoder as a query ($Q$) and introduce learnable embeddings $e_{unit}$ as keys ($K$) and values ($V$) to cross-attention based on [60]. The attention weights $\mathcal{A}_{unit}$ are computed as follows:
@@
Fig. 1e shows the process of learning attention weights to predict units and the modeling of the duration of the predicted units. We introduce an additional loss term to guide the attention module to learn semantic information. This approach implies a direct classification task, correlating the attention weights $\mathcal{A}_{unit}$ with the target unit sequence $y$.
@@
where $L$ is the length of the unit sequence and $C$ is the number of unit classes, $\mathcal{A}_{unit}^{i,j}$ represents the predicted probability of the $i$-th element for class $j$ from the attention weights $\mathcal{A}_{unit}$, and $y_{i,j}$ is the one-hot encoded class label for the $i$-th unit for class $j$. We adopted the HuBERT unit as our target unit. This is an intended design feature that ensures that the context reflected in the unit aligner is consistent with the target style, rather than based on the context of the input speech. During emotional conversion, the style autoencoder removes the source emotion from the input feature and applies the target emotion. The unit aligner then forms a unit-level context for the target emotional style.
@@ III-D Duration Modeling
The unit sequence can be predicted by identifying the focus of the attention module.
@@
where $\hat{y}^{i}$ is the $i$-th predicted unit and $\mathcal{A}_{unit}^{i}$ is the $i$-th frame of the attention weights. To extract a distinct sequence of units and their consecutive counts, a deduplication operation is applied, represented as:
@@
where $\hat{y}_{uniq}$ and $n_{count}$ denote unique units and their consecutive counts, respectively. For example, given an input sequence $\hat{y}=[4,4,2,2,2,2,1,1]$, the deduplication operation results in $\hat{y}_{uniq}=[4,2,1]$ and $n_{count}=[2,4,2]$. This implies two units with index $4$, followed by four units with index $2$ and two units with index $1$. We train the duration predictor with $n_{count}$ as the target duration. This duration is used to perform unit-level pooling to semantically bundle the output of the unit aligner. Fig. 2a explains unit-level pooling. The output of the unit aligner, $z_{attn}$, is averaged based on the duration of the unit and results in downsampling the sequence length for alignment at the unit level.
@@ III-E Hierarchical Stylize Encoder
The hierarchical stylize encoder [61] functions at two levels: unit level and frame level. It consists of two components: the unit-level stylize transformer (UST) and the frame-level stylize transformer (FST). UST processes $z_{u}$ into a $z_{us}$, denoted as $z_{us}=\text{UST}(z_{u},w_{tgt})$, focusing on unit-specific features. This refined variable $z_{u}$ is scaled to the frame level through a length regulator $\mathcal{LR}$, depicted in Fig. 2b.
@@
The duration predictor takes $z_{us}$ and $w_{tgt}$ as input and is trained to predict the unit-level duration $n_{count}$. For emotion-based duration dynamics, we introduce the flow-based stochastic duration predictor proposed in [62] to introduce duration uncertainty. The duration predictor training objective $\mathcal{L}_{dur}$ follows a negative variational lower bound.
@@ IV-A Experimental Setup
We conducted experiments using the emotional speech dataset (ESD) [^1] https://github.com/HLTSingapore/Emotional-Speech-Data[64], which contains 350 parallel utterances spoken by 10 native Mandarin speakers and 10 English speakers with 5 emotional states (neutral, happy, angry, sad, and surprise). Following the data partitioning guidelines provided by ESD, we constructed the training set with 300 samples per emotion per speaker, for a total of 15,000 samples. The validation set included 20 samples for each emotion per speaker, totaling 1,000 samples, and the test set comprised 30 samples for each emotion per speaker, totaling 1,500 samples. We sampled audio at 16,000 Hz and transformed it to an 80-bin Mel-spectrogram using a short-time Fourier transform (STFT) with a window length of 1,024 and a hop size of 256.
@@ IV-B Implementation Details
In our experimental setup, we configured both the de-stylize and stylize transformers with specific parameters: the hidden dimension, kernel size, number of heads, FFN kernel size, and feed forward network (FFN) hidden size were set to 256, 5, 2, 9, and 1024, respectively. The $\alpha$ parameter of the Beta distribution for MixLN was fixed at 0.1. All transformers used in our model were organized into $N$ layers, with $N$ established at 4. The unit aligner featured multi-head attention with 16 heads. We set $T=1$, $\beta_{t}=\beta_{0}+(\beta_{1}-\beta_{0})t$, $\beta_{0}=0.05$, and $\beta_{1}=20$ as noise schedules. The U-Net in our model was set to downsample four times and had a hidden dimension of 128. We set the inference timestep to 100. The duration predictor, which comprises residual blocks using dilated and depth-separable convolution, was structured in four layers. To address the resolution disparity between the HuBERT unit and the Mel-spectrogram, we expanded the hidden representation with a length regulator and employed linear interpolation for upsampling. In training the generator, we utilized random segments, setting the segment size to 32 frames of the Mel-spectrogram. The AdamW optimizer was used, with a learning rate of $1\times 10^{-4}$. We set the batch size to 16 and the training steps to 500K. We trained the vocoder using the official BigVGAN [^2] https://github.com/NVIDIA/BigVGAN[65] implementation, incorporating LibriTTS [66], VCTK  [^3] https://datashare.ed.ac.uk/handle/10283/2651, and ESD datasets. All comparison models were trained using a single NVIDIA RTX A6000 GPU. For broader accessibility, the code [^4] https://github.com/hs-oh-prml/DurFlexEVC and a demo [^5] https://prml-lab-speech-team.github.io/durflex/ of our proposed method are available online.
@@ IV-C1 Subjective Metrics
We conducted subjective evaluations using Amazon Mechanical Turk (mTurk). Our analysis included the mean opinion score (MOS) for naturalness (nMOS) and speaker similarity (sMOS), using a 9-point scale, ranging from 1 to 5, with increments of 0.5 units. The results are presented with a confidence interval (CI) of 95%. Furthermore, we use the emotion mean opinion classification (eMOC) as suggested in [42].
@@ IV-C2 Objective Metrics
For our objective evaluation, we incorporate a range of metrics: predicted mean opinion score, phoneme error rate (PER), character error rate (CER), word error rate (WER), emotion classification accuracy (ECA), and speaker embedding cosine similarity (SECS). The predicted MOS was assessed using UTMOS [67] [^6] https://github.com/tarepan/SpeechMOS. For the PER calculation, we used a wav2vec2.0-based phoneme recognition model from Hugging Face [68]. CER and WER were determined using Whisper [^7] https://github.com/openai/whisper[69]. In assessing SECS, we extracted speaker embeddings from both target and generated audio using Resemblyzer [^8] https://github.com/resemble-ai/Resemblyzer, subsequently computing their cosine similarity. This similarity measure ranges from -1 to 1, where higher values denote greater similarity. We evaluated the similarity for samples that shared the same speaker and emotion and then averaged these across all speakers. The objective evaluation of emotions in the generated speech was conducted using a pre-trained speech emotion recognition (SER) model. To measure SER accuracy, we employed emotion2vec [70]. We used emotion2vec+ base [^9] https://github.com/ddlBoJack/emotion2vec, a pre-trained model that supports nine classes, and only used the five sentiment classes in the ESD dataset for evaluation. We propose the emotion embedding cosine similarity (EECS) to evaluate the emotion of synthesized speech. The EECS is obtained by computing the cosine similarity of the emotion embedding between the synthesized audio and arbitrary reference audio with the target emotion. The emotion embedding was obtained using emotion2vec. Fig. 3 is a visualization of the features in emotion2vec, which shows that it encodes emotions independently of the speaker. We also evaluated the root mean square error (RMSE) for pitch and energy and calculated the difference of duration (DDUR) to assess prosody. The pitch was extracted using parselmouth [^10] https://parselmouth.readthedocs.io/en/stable/ in Hz, and the energy was obtained as the L2-norm of the absolute value of the linear spectrogram. DDUR was obtained in the same way as in [22]. Table I presents the results of each pre-trained baseline model on the test set, comprising 1500 samples, for both ground-truth and vocoded samples.
@@ IV-D Comparison Models
  • Seq2seq-EVC [^12] https://github.com/KunZhou9646/seq2seq-EVC[21]: Employing a sequence-to-sequence (seq2seq) framework, this model adopts a two-stage strategy utilizing the TTS model. A notable feature of Seq2seq-EVC is its ability to jointly model duration and pitch.
@@
  • Textless-EVC [^15] https://github.com/facebookresearch/fairseq/tree/main/examples/emotion_conversion[42]: This model approaches speech synthesis by deconstructing the speech signal into discrete learned representations. These include speech content units, prosodic features, speaker identity, and emotions. Each element is modified to align with the target emotion before being synthesized back into speech.
@@
Model UTMOS PER CER WER ECA EECS SECS
StarGAN-EVC 1.47 70.83 44.49 67.71 39.5 0.34 0.61
Seq2seq-EVC 1.54 37.29 21.68 36.87 40.0 0.39 0.63
Emovox 2.05 29.25 17.18 31.37 49.33 0.48 0.68
Mixed Emotion 2.02 29.86 18.21 33.09 57.75 0.55 0.67
Textless-EVC 2.37 22.88 12.49 23.98 56.18 0.58 0.68
DurFlex-EVC 3.39 17.31 8.26 20.75 88.64 0.85 0.75
@@ V-A Comparison of evaluation results for baseline models
To evaluate the performance of the proposed model, we conducted both objective and subjective evaluations to compare it with the baseline models. Table II shows the results of subjective evaluation. Table III shows the objective evaluation results. As the nMOS results show, the naturalness of the speech generated by our method outperforms other models. In terms of speaker similarity, our approach scored the highest on both sMOS and SECS. Fig. 4 shows the resulting SECS for all combinations of emotional conversions. This shows that the proposed model is more robust in terms of speaker similarity than the comparison models for all combinations of transformations. This means that our model is more robust in terms of speaker similarity than the comparison models for all combinations of transformations. Furthermore, the results of eMOC, ECA and EECS demonstrate that our method performs better in terms of perceptual quality as well as objective metrics. Fig. 5 shows the resulting EECS for all combinations of emotion transformations. This indicates that the proposed model synthesized speech with a higher emotional similarity than the comparison models for all combinations of transformations. The ASR evaluation also shows that our method achieves lower values in PER, CER, and WER compared to other models. This emphasizes the ability of our method to synthesize precisely pronounced speech. For a quantitative assessment of prosody, we also compared pitch and energy duration. Table LABEL:prosody shows the evaluation results for each emotion. We found that the proposed model scored better than the other models in all prosody evaluations.
TABLE IV: Comparison of prosody features (pitch, energy, duration)
Pitch

### v4

@@ III-C Unit Aligner
The unit aligner is responsible for predicting unit sequences by modeling the semantic context of the input speech. As illustrated in Fig. 1d, the unit aligner leverages a cross-attention mechanism combined with learnable embeddings to generate attention weights that guide unit prediction.
We use the output of the style autoencoder as a query ($Q$) and introduce learnable embeddings $e_{unit}$ as keys ($K$) and values ($V$) to cross-attention. The attention weights $\mathcal{A}_{unit}$ are computed as follows:
@@
Fig. 1e shows the process of learning attention weights to predict units and the modeling of the duration of the predicted units. We introduce an additional loss term to guide the attention module to learn semantic information. This approach implies a direct classification task by aligning the attention weights $\mathcal{A}_{unit}$ with the target unit sequence $y$:
@@
where $L$ is the length of the unit sequence and $C$ is the number of unit classes, $\mathcal{A}_{unit}^{i,j}$ represents the predicted probability of the $i$-th element for class $j$ from the attention weights $\mathcal{A}_{unit}$, and $y_{i,j}$ is the one-hot encoded class label for the $i$-th unit for class $j$. We adopt the HuBERT unit as our target unit. This is an intended design feature that ensures that the context reflected in the unit aligner is consistent with the target style, rather than based on the context of the input speech. During the emotional voice conversion process, the style autoencoder first removes the source emotion from the input features and applies the target emotion. The unit aligner then processes these stylized features to predict the corresponding unit sequences. This design enables the model to transform context efficiently. It also supports parallel processing without relying on seq2seq structures.
@@ III-D Duration Modeling
Once the unit aligner has predicted the unit sequence, the duration modeling component determines the duration of each unit based on consecutive identical units. This allows the model to adjust the generated speech length to align with the target emotional style. The unit sequence $\hat{y}$ is obtained by selecting the unit with the highest probability for each frame:
@@
where $\hat{y}^{i}$ is the $i$-th predicted unit and $\mathcal{A}_{unit}^{i}$ represents the attention weights for the $i$-th frame. To extract a distinct sequence of units and their consecutive counts, a deduplication operation is applied:
@@
where $\hat{y}_{uniq}$ contains the sequence of unique units, and $n_{count}$ represents the number of consecutive occurrences for each unit. For example, given an input sequence $\hat{y}=[4,4,2,2,2,2,1,1]$, the deduplication results in $\hat{y}_{uniq}=[4,2,1]$ and $n_{count}=[2,4,2]$. This indicates two instances of unit $4$, followed by four instances of unit $2$ and two instances of unit $1$.
The duration predictor is trained using $n_{count}$ as the target durations. This predictor estimates the number of consecutive units based on the unit-level representations generated by the unit aligner. Fig. 2a explains unit-level pooling. The output of the unit aligner, $z_{attn}$, is averaged based on the duration of the unit and results in downsampling the sequence length for alignment at the unit level.
@@ III-E Hierarchical Stylize Encoder
The hierarchical stylize encoder [57] functions at two levels: the unit level and the frame level. It consists of two components: the unit-level stylize transformer (UST) and the frame-level stylize transformer (FST). UST processes $z_{u}$ into a $z_{us}$, denoted as $z_{us}=\text{UST}(z_{u},w_{tgt})$, focusing on unit-specific features. This refined variable $z_{u}$ is scaled to the frame level through a length regulator $\mathcal{LR}$, depicted in Fig. 2b.
@@
The duration predictor takes $z_{us}$ and $w_{tgt}$ as input and is trained to predict the unit-level duration $n_{count}$. For emotion-based duration dynamics, we introduce the flow-based stochastic duration predictor proposed in [58] to add duration uncertainty. The duration predictor training objective $\mathcal{L}_{dur}$ follows a negative variational lower bound.
@@ IV-A Experimental Setup
We conducted experiments using the emotional speech dataset (ESD) [^1] https://github.com/HLTSingapore/Emotional-Speech-Data[59], which contains 350 parallel utterances spoken by 10 native Mandarin speakers and 10 English speakers with 5 emotional states (neutral, happy, angry, sad, and surprise). Following the data partitioning guidelines provided by ESD, we constructed the training set with 300 samples per emotion per speaker, for a total of 15,000 samples. The validation set included 20 samples for each emotion per speaker, totaling 1,000 samples, and the test set comprised 30 samples for each emotion per speaker, totaling 1,500 samples. Audio was sampled at 16,000 Hz and transformed into an 80-bin Mel-spectrogram using a short-time Fourier transform (STFT) with a window length of 1,024 and a hop size of 256.
@@ IV-B Implementation Details
In our experimental setup, we configured both the de-stylize and stylize transformers with specific parameters: the hidden dimension, kernel size, number of heads, FFN kernel size, and feed forward network (FFN) hidden size were set to 256, 5, 2, 9, and 1024, respectively. The $\alpha$ parameter of the Beta distribution for MixLN was fixed at 0.1. All transformers used in our model were organized into $N$ layers, with $N$ set to 4. The unit aligner featured multi-head attention with 16 heads. We set $T=1$, $\beta_{t}=\beta_{0}+(\beta_{1}-\beta_{0})t$, $\beta_{0}=0.05$, and $\beta_{1}=20$ as noise schedules. The U-Net in our model was set to downsample four times and had a hidden dimension of 128. We evaluated our model with two different time steps: 4 and 100. The duration predictor, which comprises residual blocks using dilated and depth-separable convolution, was structured in four layers. To address the resolution disparity between the HuBERT unit and the Mel-spectrogram, we expanded the hidden representation with a length regulator and employed linear interpolation for upsampling. In training the generator, we utilized random segments, setting the segment size to 32 frames of the Mel-spectrogram. The AdamW optimizer was used, with a learning rate of $1\times 10^{-4}$. We set the batch size to 16 and the number of training steps to 500K. We trained the vocoder using the official BigVGAN [^2] https://github.com/NVIDIA/BigVGAN[60] implementation, incorporating the LibriTTS [61], VCTK  [^3] https://datashare.ed.ac.uk/handle/10283/2651, and ESD datasets. All experiments were conducted on an Intel® Xeon® Gold 6148 CPU @ 2.40 GHz and a single NVIDIA RTX A6000 GPU. For broader accessibility, the code [^4] https://github.com/hs-oh-prml/DurFlexEVC and a demo [^5] https://prml-lab-speech-team.github.io/durflex/ of our proposed method are available online.
@@ IV-C1 Subjective Metrics
We conducted subjective evaluations using Amazon Mechanical Turk (mTurk). Our analysis included the mean opinion score (MOS) for naturalness (nMOS) and speaker similarity (sMOS), using a 9-point scale ranging from 1 to 5, with increments of 0.5 units. The results are presented with a confidence interval (CI) of 95%. Furthermore, we use the emotion mean opinion classification (eMOC) as suggested in [27].
@@ IV-C2 Objective Metrics
For our objective evaluation, we incorporate a range of metrics: predicted mean opinion score, phoneme error rate (PER), character error rate (CER), word error rate (WER), emotion classification accuracy (ECA), and speaker embedding cosine similarity (SECS). The predicted MOS was assessed using UTMOS [62] [^6] https://github.com/tarepan/SpeechMOS. For the PER calculation, we used a wav2vec2.0-based phoneme recognition model from Hugging Face [63]. CER and WER were determined using Whisper [^7] https://github.com/openai/whisper[64]. In assessing SECS, we extracted speaker embeddings from both target and generated audio using Resemblyzer [^8] https://github.com/resemble-ai/Resemblyzer, subsequently computing their cosine similarity. This similarity measure ranges from -1 to 1, where higher values denote greater similarity. We evaluated the similarity for samples that shared the same speaker and emotion and then averaged these across all speakers. The objective evaluation of emotions in the generated speech was conducted using a pre-trained speech emotion recognition (SER) model.
To measure SER accuracy, we employed emotion2vec [65]. We used emotion2vec+ base [^9] https://github.com/ddlBoJack/emotion2vec, a pre-trained model that supports nine classes, and only used the five sentiment classes in the ESD dataset for evaluation. We propose the emotion embedding cosine similarity (EECS) to evaluate the emotion of synthesized speech. The EECS is obtained by computing the cosine similarity of the emotion embedding between the synthesized audio and arbitrary reference audio with the target emotion. The emotion embedding was obtained using emotion2vec. Fig. 3 is a visualization of the features in emotion2vec, which shows that it encodes emotions independently of the speaker. We also evaluated the root mean square error (RMSE) for pitch and energy and calculated the difference of duration (DDUR) to assess prosody. The pitch was extracted using parselmouth [^10] https://parselmouth.readthedocs.io/en/stable/ in Hz, and the energy was obtained as the L2-norm of the absolute value of the linear spectrogram. DDUR was obtained in the same way as in [22]. We measured the real-time factor (RTF) of inference time for computational analysis. Table I presents the results of each pre-trained baseline model on the test set, comprising 1500 samples, for both ground-truth and vocoded samples.
@@ IV-D Comparison Models
  • Seq2seq-EVC [^12] https://github.com/KunZhou9646/seq2seq-EVC[21]: Employing a seq2seq framework, this model adopts a two-stage strategy utilizing the TTS model. A notable feature of Seq2seq-EVC is its ability to jointly model duration and pitch.
@@
  • Textless-EVC [^15] https://github.com/facebookresearch/fairseq/tree/main/examples/emotion_conversion[27]: This model approaches speech synthesis by deconstructing the speech signal into discrete learned representations. These include speech content units, prosody features, speaker identity, and emotions. Each element is modified to align with the target emotion before being synthesized back into speech.
@@
Model UTMOS PER CER WER ECA EECS SECS RTF
StarGAN-EVC 1.47 70.83 44.49 67.71 39.5 0.34 0.61 0.0051
Seq2seq-EVC 1.54 37.29 21.68 36.87 40.0 0.39 0.63 0.2325
Emovox 2.05 29.25 17.18 31.37 49.33 0.48 0.68 0.2109
Mixed Emotion 2.02 29.86 18.21 33.09 57.75 0.55 0.67 0.2289
Textless-EVC 2.37 22.88 12.49 23.98 56.18 0.58 0.68 0.4340
DurFlex-EVC (4 time steps) 3.21 18.89 8.61 21.36 89.52 0.85 0.71 0.1334
DurFlex-EVC (100 time steps) 3.39 17.31 8.26 20.75 88.64 0.85 0.75 1.9200
@@ V-A Comparison of evaluation results for baseline models
To evaluate the performance of the proposed model, we conducted both objective and subjective evaluations to compare it with the baseline models. Table II shows the results of subjective evaluation. Table III shows the objective evaluation results. As the nMOS results show, the naturalness of the speech generated by our method outperforms that of other models. In terms of speaker similarity, our approach scored the highest on both sMOS and SECS. Fig. 4 shows the resulting SECS for all combinations of emotional conversions. This demonstrates that the proposed model is more robust in terms of speaker similarity than the comparison models for all combinations of transformations. Furthermore, the results of eMOC, ECA and EECS demonstrate that our method performs better in terms of perceptual quality as well as objective metrics. Fig. 5 shows the resulting EECS for all combinations of emotion transformations. This indicates that the proposed model synthesized speech with a higher emotional similarity than the comparison models for all combinations of transformations. The ASR evaluation also shows that our method achieves lower values in PER, CER, and WER compared to other models. This emphasizes the ability of our method to synthesize precisely pronounced speech.
We evaluated the computational efficiency of each model by comparing their RTF. Autoregressive models, including Seq2seq-EVC, Emovox, Mixed Emotion, and Textless-EVC, generally demonstrated slower inference speeds compared to parallel generation models such as StarGAN-EVC. DurFlex-EVC, as a diffusion-based model, exhibited varying RTFs depending on the number of sampling time steps. With just four time steps, DurFlex-EVC achieved an RTF of 0.1334, significantly faster than all compared autoregressive models. At this setting, DurFlex-EVC maintained competitive performance and exceeded the speed and performance metrics of other autoregressive models. However, when using 100 time steps, DurFlex-EVC achieved the highest performance, though at a slower RTF of 1.9200 due to the increased computational demand of the diffusion process.
For a quantitative assessment of prosody, we also compared pitch and energy duration. Table LABEL:prosody shows the evaluation results for each emotion. We found that the proposed model scored better than the other models in all prosody evaluations.
V-B Experiments for Analyzing Model Architectures
V-B1 Effectiveness of Style Autoencoder
We conducted an experiment to verify the effectiveness of the style autoencoder (SAE). w/o SAE is a model that stacks a standard feed forward transformer block with a full layer of SAE, without perturbation and conditioning for emotion style. In Table V, w/o SAE shows the results of the ablation study on it. The overall performance degradation observed in models without SAE indicates the importance of disentangling content and style from input features. Fig. 6 shows the change in cross entropy for styles as input features pass through SAE. To obtain cross entropy, we trained a classifier for emotion style on features extracted from all layers of SAE. The black line is the result for the model with SAE and the orange line is the result for the model without SAE. The w/o SAE model shows high cross entropy across the layers, while the model with SAE shows increasing cross entropy as it passes through the de-stylize transformer layers and decreasing cross entropy as it passes through the stylize transformer layers. This means that the style is disentangled and conditioned from the feature by SAE.
We also experimented with adversarial training strategies to remove source styles from input features. This replaces the de-stylize transformer in the style autoencoder with a standard transformer, and adds a gradient reversal layer and classifier to prevent its output from learning the style. This adversarial training strategy is often used to disentangle specific styles [67, 68, 69].
The results for this are shown in Table V as w/ adv, and show a performance degradation on all metrics except EECS.
TABLE IV: Comparison of prosody features (pitch, energy, duration).
Pitch (RMSE)
@@ V-B1 Effectiveness of Style Autoencoder
V-B2 Effectiveness of Unit Aligner
The unit aligner (UA) was introduced to model stylized contexts. Table V w/o UA shows the results of the ablation experiment for this. The results showed that w/o UA performed better in terms of voice quality and pronunciation but not in terms of emotion conversion. This indicates that there remains a lot of information about the source style in the features and that the UA functions as a bottleneck and has a significant impact on style control. We measured the Jensen-Shannon divergence (JSD) to validate that the predicted units in UA reflect the target emotion. Table VI shows the JSD results of the units for each emotion in Textless-EVC and DurFlex-EVC. DurFlex-EVC achieved a better JSD than Textless-EVC for all emotions, indicating that it provides an appropriate representation of the context.
V-B3 Effectiveness of Hierarchical Stylize Encoder
The output of the unit aligner is a representation in units, which is expanded to frame-level by a duration predictor and length regulator. For the frame-level stylization in this representation, we introduce a hierarchical stylize encoder (HSE). The results of the ablation study without the HSE are shown in Table V under the entry labeled w/o HSE. The evaluation results showed an overall performance degradation with w/o HSE. This is inferred to be due to HSE reducing the burden on the Mel-spectrogram generator, resulting in improved generation quality.
(a) Angry
(b) Happy
(c) Sad
(d) Surprise